# Searching for blue in the dark

### v1

@@ Abstract
The primordial power spectrum of curvature perturbations has been well-measured on large scales but remains fairly unconstrained at smaller scales, where significant deviations from $\Lambda$CDM may occur. Measurements of 21-cm intensity mapping in the dark ages promise to access very small scales that have yet to be probed, extending beyond the reach of CMB and galaxy surveys. In this paper, we investigate how small-scale power-law enhancements—or blue tilts—of the primordial power spectrum affect the 21-cm power spectrum. We consider generic enhancements due to curvature modes, isocurvature modes, and runnings of the spectral tilt. We present forecasts for Earth- and lunar-based instruments to detect a blue-tilted primordial spectrum. We find that an Earth-based instrument capable of reaching the dark ages could detect any enhancements of power on nearly all the scales it can observe, which depends on the baseline of the interferometer. The smallest scales observed by such an instrument can only detect a very strong enhancement. However, an instrument on the far side of the Moon of the same size would be able to probe shallower slopes with higher precision. We forecast results for instruments with $100\,{\rm km}\,(3000\,{\rm km})$ baselines and find that they can probe up to scales of order $k_{\rm max}\sim 8\,{\rm Mpc}^{-1}\,(k_{\rm max}\sim 250\,{\rm Mpc}^{-1})$, thereby providing invaluable information on exotic physics and testing inflationary models on scales not otherwise accessible.
@@ I Introduction
The power spectrum of primordial perturbations describes the initial conditions of the Universe and thus of structure formation. It is well-characterized by curvature perturbations $\zeta$ on relatively scales from roughly $10^{-4}$ to $1\,{\rm Mpc}^{-1}$, by measurements of cosmic microwave background (CMB) anisotropies and large-scale structure (LSS) Aghanim et al. (2020); Kalaja et al. (2019); Ivanov et al. (2020); Philcox and Ivanov (2022); Ivanov (2021); Collaboration (2024); Akrami et al. (2020).
@@
Previous work has constrained the primordial power spectrum on smaller scales using the abundance of ultracompact minihalos Bringmann et al. (2012), the distribution of stars in an ultra-faint dwarf galaxy Graham and Ramani (2024), spectral distortions of the CMB Silk (1981); Chluba et al. (2021), and the high-redshift UV galaxy luminosity function Sabti et al. (2022); Yoshiura et al. (2020). However, these methods can have large astrophysical uncertainties, preventing them from achieving the same level of sensitivity and robustness as the constraints from CMB anisotropies and LSS. Possible future CMB experiments, such as PIXIE Kogut et al. (2024) and CMB-HD Collaboration (2022), could reach scales up to $k\sim 10\,{\rm Mpc}^{-1}$.
@@
Probing the primordial power spectrum on small scales could be the key to discovering new physics Bechtol et al. (2022); Boddy et al. (2022). Many exotic physics models generate an increase (or blue tilt) of the small-scale power spectrum, such as nonstandard inflationary scenarios Martin and Brandenberger (2001); Gong and Sasaki (2011); Germani and Prokopec (2017); Ezquiaga et al. (2018); Ballesteros et al. (2020); Mishra and Sahni (2020); Braglia et al. (2020); Wang et al. (2024); Mollerach et al. (1994); García-Bellido and Linde (1997); Barrow and Liddle (1993); Palma et al. (2020), gravitational particle production of dark matter Chung et al. (2005); Chung and Yoo (2013); Chung et al. (2015); Ling and Long (2021); Kolb and Long (2023); Graham et al. (2016), axions Chung and Upadhye (2017); Chung and Tadepalli (2024), quantum decoherence during inflation Martin and Vennin (2018), and bouncing cosmologies Qiu et al. (2011); Cai et al. (2012). Such blue-tilted primordial (or matter) power spectra have been investigated using CMB/LSS Chung and Upadhye (2018), spectral distortions Chluba and Grin (2013), dark matter substructure Ando et al. (2022); Graham and Ramani (2024); Josan and Green (2010); Bringmann et al. (2012), quasar light curves Karami et al. (2018), Lyman-$\alpha$ forest Afshordi et al. (2003), and Milky Way satellite velocities Esteban et al. (2023); Dekker and Kravtsov (2024). There have also been forecasts for constraining blue tilts with Euclid and MegaMapper Chung et al. (2023), as well as for SKAO using 21-cm line-intensity-mapping (LIM) at relatively low redshifts $(z<20)$ Sekiguchi et al. (2014).
@@
In this work, we consider the sensitivity that future 21-cm fluctuation measurements from the dark ages have to enhancements of small-scale power, with respect to $\Lambda$CDM. The main benefit of measuring the 21-cm power spectrum during the dark ages Pritchard and Loeb (2012); Furlanetto et al. (2006); Zaldarriaga et al. (2004a); Loeb and Zaldarriaga (2004a); Zaldarriaga et al. (2004b); Loeb and Zaldarriaga (2004b); Furlanetto (2006), which spans the redshift range $30\lesssim z\lesssim 200$, is that there is no astrophysical contamination, since the first stars have not yet formed; thus, we can more directly probe early Universe cosmology. Moreover, fluctuations are not affected by Silk damping Silk (1968) and therefore remain undamped down to the baryon Jeans scale $k_{J}\sim 300\,{\rm Mpc}^{-1}$. As a result, we can access modes at much smaller scales with 21-cm LIM, compared to the CMB.
@@
There are several proposals for low-frequency interferometers that would be able to detect the dark ages 21-cm signal. Most are lunar-based experiments Burns et al. (2017); Burns et al. (2019a); Burns et al. (2019b); Bale et al. (2023); Chen et al. (2023); Polidan et al. (2024); Lazio et al. (2009); Price et al. (2018), as well as a lunar-orbiting CubeSat Artuc and Acedo (2024). It was also the aim of several Earth-based instruments to observe the dark ages Bevins et al. (2022); Jishnu Nambissan et al. (2021); Blyth et al. (2015); Maio et al. (2015); Monsalve et al. (2024). While the current designs are not yet capable of reaching such high redshifts, future versions of them potentially could. We consider the sensitivity to enhancements of small-scale power for a few hypothetical instruments, both Earth- and lunar-based.
@@
Previous studies have investigated how changes to the matter power spectrum affect the 21-cm signal Jones et al. (2021); Vanzan et al. (2024); Hotinli et al. (2022); Flitter and Kovetz (2022); Muñoz et al. (2020, 2017); Cole and Silk (2020); Short et al. (2022); Driskell et al. (2022); Ali-Haïmoud et al. (2014); most of them consider a suppression or enhancement of small-scale structure for particular models. We take an agnostic approach by considering generic power-law enhancements of the primordial power spectrum due to a small-scale enhancement of curvature modes, the addition of uncorrelated cold dark matter isocurvature modes to the standard $\Lambda$CDM curvature modes, and a running of the spectral tilt. We find that an Earth-based instrument can already probe most power spectra increases on the observable scales, while a lunar-based instrument of the same size can probe shallower slopes and obtain a higher accuracy. When increasing the size of the instrument on the Moon, one can probe a much larger range of scales, allowing revolutionary and stringent tests of exotic physics and inflationary models.
@@
This paper is organized as follows. In Section II we introduce different blue-tilted power spectra considered in this work. In Section III we give a brief overview of 21-cm LIM from the dark ages. In Section IV we describe our methods for measuring the blue-tilted power spectrum and provide the instrument specifications we use for our forecasts. In Section V we show the results of our forecasts. We summarize our findings and conclude in Section VI.
@@ II Blue-tilted primordial power spectrum
is characterized by a tilt $n_{s}$ and amplitude $A_{s}$ at the pivot scale $k_{0}$. Measurements of CMB anisotropy from the Planck satellite, in conjunction with LSS data, show that this power spectrum is nearly scale-invariant and red-tilted (i.e., decreases with $k$) with $n_{s}=0.9649\pm 0.0084$ ($95\%$ CL, Planck TT, TE, EE + lowE + lensing), and has an amplitude given by $\ln(10^{10}A_{s})=3.044\pm 0.028$ ($95\%$ CL, Planck TT, TE, EE + lowE + lensing) at $k_{0}=0.05\,{\rm Mpc}^{-1}$ Aghanim et al. (2020). These values are compatible with the simplest prediction of single-field, slow-roll inflation.
@@
Small-scale enhancements arise from a variety of early Universe physics, such as inflationary scenarios Martin and Brandenberger (2001); Gong and Sasaki (2011); Germani and Prokopec (2017); Ezquiaga et al. (2018); Ballesteros et al. (2020); Mishra and Sahni (2020); Braglia et al. (2020); Wang et al. (2024); Mollerach et al. (1994); García-Bellido and Linde (1997); Barrow and Liddle (1993); Palma et al. (2020), gravitational particle production of dark matter Chung et al. (2005); Chung and Yoo (2013); Chung et al. (2015); Ling and Long (2021); Kolb and Long (2023); Graham et al. (2016), axions Chung and Upadhye (2017); Chung and Tadepalli (2024), quantum decoherence during inflation Martin and Vennin (2018), and bouncing cosmologies Qiu et al. (2011); Cai et al. (2012). These scenarios produce blue-tilted enhancements in the total primordial power spectrum $\Delta^{2}(k)$, either by directly increasing curvature power spectrum at small scales or by generating small-scale cold dark matter isocurvature.
@@ II.1 Enhanced curvature power spectrum
We start by considering primordial blue-tilted curvature power spectra at small scales, which can be produced by e.g. Mollerach et al. (1994); Martin and Brandenberger (2001); Gong and Sasaki (2011); Germani and Prokopec (2017); Ezquiaga et al. (2018); Ballesteros et al. (2020); Mishra and Sahni (2020); Wang et al. (2024); García-Bellido and Linde (1997); Barrow and Liddle (1993); Palma et al. (2020); Martin and Vennin (2018); Qiu et al. (2011). The primordial spectrum may possess additional features, such as a turnover that produces a bump at some large value of $k$. Since such features are model-dependent, we do not account for them. However, we impose a threshold of $\Delta_{\zeta}^{2}(k)\leq 0.01$, forcing the power spectrum to remain constant at arbitrarily large $k$. The presence of the threshold has no significant impact on our analyses. [^1] There are cases in which the threshold is imposed at scales relevant to our analyses. However, these situations occur in the signal-dominated regime, so our projections based on signal-to-noise ratios remain insensitive to the threshold.
To investigate small-scale enhancements in a model-independent manner, we adopt a simple broken-power-law parameterization of the dimensionless primordial curvature power spectrum:
@@
The power-law scaling for $k>k_{b}$ in Eq. (2) can be mapped to various models. For example, for tilted hybrid inflation García-Bellido and Linde (1997), the spectral index and the amplitude of the power spectrum can be mapped to the effective mass of the inflaton field and the value of the field at the moment when perturbations that have momentum $k$ at the end of inflation were produced. Another example is quantum decoherence during inflation Martin and Vennin (2018). The evolution of inflationary curvature perturbations can be modified due to the interaction with other degrees of freedom present in the early Universe. This interaction leads to quantum decoherence, and in various cases to a blue-tilted correction of the power spectrum on scales considered in this work. The time dependence and strength of the interaction between the perturbations and other degrees of freedom determine $n_{b}$ and $k_{b}$.
@@ II.2 Enhanced dark matter isocurvature power spectrum
A different class of small-scale enhancements arises from the production of isocurvature modes Bartolo et al. (2001); Linde and Mukhanov (1997); Takeuchi and Chongchitnan (2014); Ghosh et al. (2022); Minoda et al. (2022); Chung and Upadhye (2017); Sekiguchi et al. (2014); Kawasaki et al. (2011); Kasuya and Kawasaki (2009); Braglia et al. (2020); Chung et al. (2005, 2015); Chung and Upadhye (2017); Graham et al. (2016); Ling and Long (2021); Kolb and Long (2023); Kasuya et al. (1997); Chung and Yoo (2013); Chung and Tadepalli (2024); Wands et al. (2002). We assume that adiabatic modes give rise to the dimensionless primordial curvature power spectrum in Eq. (1), and the isocurvature modes produce a dimensionless primordial isocurvature power spectrum
@@
as demonstrated in Fig. 2. For all scenarios considered in this work, we have checked explicitly that the amplitudes are consistent with constraints from Planck Akrami et al. (2020). Similar to the curvature case, we impose a threshold of $\Delta^{2}(k)\leq 0.01$.
@@
A blue-tilted isocurvature power spectrum can arise, for example, from the QCD axion Chung and Tadepalli (2024); Kasuya and Kawasaki (2009); Chung and Tadepalli (2022); Kasuya et al. (1997), which produces an isocurvature spectral index of $1<n_{\mathrm{iso}}\leq 4$. The bounds on the isocurvature parameters can be translated to a bound on the inflationary energy scale $H_{I}$ Hamann et al. (2009); Hertzberg et al. (2008). A similar translation can be done in the context of ultralight axions Marsh et al. (2013), for which $H_{I}$ depends on $A_{\rm iso}/A_{s}$ and the fraction of dark matter that ultralight axions constitute.
Alternatively, a blue tilt with $n_{\rm iso}=4$ can be generated by particle production mechanisms with a finite correlation length. The resulting isocurvature power spectrum vanishes on scales smaller than a cutoff scale determined by causality, and only white noise is left on these smaller scales. Examples of such mechanisms are post-inflationary dark matter production Amin and Mirbabayi (2024); Caprini et al. (2009), vector dark matter Graham et al. (2016); Gorghetto et al. (2022), and PQ-breaking during inflation with subsequent axion production Redi and Tesi (2023).
@@ II.3 Running parameters
The parameterization in Eq. (2) is meant to describe the main features of the primordial spectrum; however, if there is a transition from the decreasing to the increasing behavior as a function of $k$, this transition would be described by the first and higher order derivatives of the spectrum. We parameterize the primordial power spectrum in this case as Kosowsky and Turner (1995)
@@
are the first and second running parameters of the tilt, respectively Kosowsky and Turner (1995); Baumann et al. (2009). Planck 2018 data constrains the running parameters to be $\alpha_{s}=0.013\pm 0.012$ and $\beta_{s}=0.022\pm 0.012$ (at the pivot scale $k=0.05\,\rm{Mpc}^{-1}$ghanim et al. (2020); Cabass et al. (2016); Akrami et al. (2020).
Standard single-field slow-roll inflation predicts $\alpha_{s}\sim(1-n_{s})^{2}\sim\mathcal{O}(\epsilon^{2},\eta^{2})\sim 10^{-3}$ and $\beta_{s}\sim(1-n_{s})^{3}\sim\mathcal{O}(\epsilon^{3},\eta^{3})\sim 10^{-5}$ Adshead et al. (2011); Muñoz et al. (2017); Kosowsky and Turner (1995); Lidsey et al. (1997). Therefore, high precision constraints on the running parameters could either confirm or rule out this inflationary scenario and permit tests of other inflationary models that predict a running, such as multifield Gong (2015), and warm inflation Das and Ramos (2023).
Note that this parameterization is only valid for small deviations from the scale-invariant case and is thus not appropriate for the large enhancements studied in Section II.1. We focus on the region of $k$ in which the power spectrum begins to transition to becoming blue tilted, and thus the parameterization is valid.
@@
In this section we describe the main observable of interest: the intensity mapping signal of 21-cm neutral hydrogen. We provide a brief review of key concepts for 21-cm LIM from the dark ages, before the formation of the first stars, which occurs around $z\sim 30$ in $\Lambda$CDM. [^2] We note that enhancements of small-scale structure should lead to advanced star formation. Studying the onset of cosmic dawn is beyond the scope of this work, and we use $z=30$ as a convenient reference redshift for our forecasts. For more details, we refer the reader to Refs. Madau et al. (1997); Zaldarriaga et al. (2004a); Pritchard and Loeb (2008, 2010); Furlanetto (2019) for 21-cm LIM, Refs. Loeb and Zaldarriaga (2004a); Furlanetto et al. (2006); Lewis and Challinor (2007); Furlanetto et al. (2009); Pritchard and Loeb (2012); Iliev et al. (2002) for specific application to the dark ages, and Refs. Kovetz et al. (2017); Bernal and Kovetz (2022) for recent reviews on LIM.
@@
The observed quantity for the 21-cm signal is the brightness temperature contrast with respect to CMB Ali-Haïmoud et al. (2014),
@@
The Sobolev optical depth Sobolev (1957) is
@@
The brightness temperature is a function of the local hydrogen density and the gas temperature, and it can be parameterized as Ali-Haïmoud et al. (2014)
@@
where $\delta_{b}=\delta n_{b}/\bar{n}_{b}\simeq\delta_{H}$ up to negligible corrections, $\delta_{T_{\rm gas}}=\delta T_{\rm gas}/\bar{T}_{\rm gas}$, and $\delta_{v}=-(1+z)\partial_{r}v/H(z)$. The coefficients $c_{b,T}$ are functions of redshift only; the mean brightness temperature $\bar{T}_{21}$ is defined by setting all perturbations to zero. Following Ref. Short et al. (2020), we neglect fluctuations in the gas temperature and in the neutral hydrogen fraction to write the 21-cm brightness temperature fluctuations as
@@
By combining Eq. (11) with the continuity equation, we can directly relate the 21-cm angular power spectrum to the matter power spectrum $P_{m}(k)$ Bharadwaj and Ali (2004); Ali-Haïmoud et al. (2014); Short et al. (2020):
@@
where $X=\{\zeta,\mathcal{S}_{\rm cdm}\}$ represents the initial condition, and we assume the adiabatic and cold dark matter isocurvature modes are uncorrelated. We use CLASS Blas et al. (2011) to compute the linear growth factor $D_{X}(z)$ and the matter transfer function $T_{m,X}(k)$. In the $\Lambda$CDM case, the power spectrum is still linear for the scales and redshifts we probe. In the cases where the primordial power spectrum reaches values above $10^{-3}$ at the scales of interest, the resulting matter power spectrum could become non-linear. For our results, the majority of the sensitivity is near the $\ell_{\rm{max}}$ value (see Section V for more details), and we do not expect our conclusions to be affected by assuming the matter power spectrum is still linear; a careful investigation of this issue is left for future work.

### v2

@@ Abstract
The primordial power spectrum of curvature perturbations has been well-measured on large scales but remains fairly unconstrained at smaller scales, where significant deviations from $\Lambda$CDM may occur. Measurements of 21-cm intensity mapping in the dark ages promise to access very small scales that have yet to be probed, extending beyond the reach of cosmic microwave background and galaxy surveys. In this paper, we investigate how small-scale power-law enhancements—or blue tilts—of the primordial power spectrum affect the 21-cm power spectrum. We consider generic enhancements due to curvature modes, isocurvature modes, and runnings of the spectral tilt. We present forecasts for Earth- and lunar-based instruments to detect a blue-tilted primordial spectrum. We find that an Earth-based instrument capable of reaching the dark ages could detect any enhancements of power on nearly all the scales it can observe, which depends on the baseline of the interferometer. The smallest scales observed by such an instrument can only detect a very strong enhancement. However, an instrument on the far side of the Moon of the same size would be able to probe shallower slopes with higher precision. We forecast results for instruments with $100\,{\rm km}\,(3000\,{\rm km})$ baselines and find that they can probe up to scales of order $k_{\rm max}\sim 8\,{\rm Mpc}^{-1}\,(k_{\rm max}\sim 250\,{\rm Mpc}^{-1})$, thereby providing invaluable information on exotic physics and testing inflationary models on scales not otherwise accessible.
@@ I Introduction
The power spectrum of primordial perturbations describes the initial conditions of the Universe and thus of structure formation. It is well-characterized by curvature perturbations $\zeta$ on relative scales from roughly $10^{-4}$ to $1\,{\rm Mpc}^{-1}$, by measurements of cosmic microwave background (CMB) anisotropies and large-scale structure (LSS) [1, 2, 3, 4, 5, 6, 7].
@@
Previous work has constrained the primordial power spectrum on smaller scales using the abundance of ultracompact minihalos [8], the distribution of stars in an ultrafaint dwarf galaxy [9], spectral distortions of the CMB [10, 11], and the high-redshift UV galaxy luminosity function [12, 13]. However, these methods can have large astrophysical uncertainties, preventing them from achieving the same level of sensitivity and robustness as the constraints from CMB anisotropies and LSS. Possible future CMB experiments, such as PIXIE [14] and CMB-HD [15], could reach scales up to $k\sim 10\,{\rm Mpc}^{-1}$.
@@
Probing the primordial power spectrum on small scales could be the key to discovering new physics [16, 17]. Many exotic physics models generate an increase (or blue tilt) of the small-scale power spectrum, such as nonstandard inflationary scenarios [18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35], gravitational particle production of dark matter [36, 37, 38, 39, 40, 41, 42], axions [43, 44, 45, 46, 47], quantum decoherence during inflation [48], and bouncing cosmologies [49, 50]. Such blue-tilted primordial (or matter) power spectra have been investigated using CMB/LSS [51, 52], spectral distortions [53], dark matter substructure [54, 9, 55, 8], quasar light curves [56], Lyman-$\alpha$ forest [57, 58], and Milky Way satellite velocities [59, 60]. There have also been forecasts for constraining blue tilts with Euclid and MegaMapper [61], as well as for Square Kilometer Array Observatory (SKAO) using 21-cm line-intensity-mapping (LIM) at relatively low redshifts $(z<20)$ [62].
@@
In this work, we consider the sensitivity that future 21-cm fluctuation measurements from the dark ages have to enhancements of small-scale power, with respect to $\Lambda$CDM. The main benefit of measuring the 21-cm power spectrum during the dark ages [63, 64, 65, 66, 67, 68], which spans the redshift range $30\lesssim z\lesssim 200$, is that there is no astrophysical contamination, since the first stars have not yet formed; thus, we can more directly probe early Universe cosmology. Moreover, fluctuations are not affected by Silk damping [69] and therefore remain undamped down to the baryon Jeans scale $k_{J}\sim 300\,{\rm Mpc}^{-1}$. As a result, we can access modes at much smaller scales with 21-cm LIM, compared to the CMB.
@@
There are several proposals for low-frequency interferometers that would be able to detect the dark ages 21-cm signal. Most are lunar-based experiments [70, 71, 72, 73, 74, 75, 76, 77], as well as a lunar-orbiting CubeSat [78]. It was also the aim of several Earth-based instruments to observe the dark ages [79, 80, 81, 82, 83]. While the current designs are not yet capable of reaching such high redshifts, future versions of them potentially could. We consider the sensitivity to enhancements of small-scale power for a few hypothetical instruments, both Earth- and lunar-based.
@@
Previous studies have investigated how changes to the matter power spectrum affect the 21-cm signal [84, 85, 86, 87, 88, 89, 90, 91, 92, 93]; most of them consider a suppression or enhancement of small-scale structure for particular models. We take an agnostic approach by considering generic power-law enhancements of the primordial power spectrum due to a small-scale enhancement of curvature modes, the addition of uncorrelated cold dark matter isocurvature modes to the standard $\Lambda$CDM curvature modes, and a running of the spectral tilt. We find that an Earth-based instrument can already probe most power spectra increases on the observable scales, while a lunar-based instrument of the same size can probe shallower slopes and obtain a higher accuracy. When increasing the size of the instrument on the Moon, one can probe a much larger range of scales, allowing revolutionary and stringent tests of exotic physics and inflationary models.
@@
This paper is organized as follows. In Sec. II we introduce different blue-tilted power spectra considered in this work. In Sec. III we give a brief overview of 21-cm LIM from the dark ages. In Sec. IV we describe our methods for measuring the blue-tilted power spectrum and provide the instrument specifications we use for our forecasts. In Sec. V we show the results of our forecasts. We summarize our findings and conclude in Sec. VI.
@@ II Blue-tilted primordial power spectrum
is characterized by a tilt $n_{s}$ and amplitude $A_{s}$ at the pivot scale $k_{0}$. Measurements of CMB anisotropy from the Planck satellite, in conjunction with LSS data, show that this power spectrum is nearly scale invariant and red tilted (i.e., decreases with $k$) with $n_{s}=0.9649\pm 0.0084$ ($95\%$ CL, Planck TT, TE, EE + lowE + lensing), and has an amplitude given by $\ln(10^{10}A_{s})=3.044\pm 0.028$ ($95\%$ CL, Planck TT, TE, EE + lowE + lensing) at $k_{0}=0.05\,{\rm Mpc}^{-1}$ [1]. These values are compatible with the simplest prediction of single-field, slow-roll inflation.
@@
Small-scale enhancements arise from a variety of early Universe physics, such as inflationary scenarios [18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29], gravitational particle production of dark matter [36, 37, 38, 39, 40, 41], axions [43, 44], quantum decoherence during inflation [48], and bouncing cosmologies [49, 50]. These scenarios produce blue-tilted enhancements in the total primordial power spectrum $\Delta^{2}(k)$, either by directly increasing curvature power spectrum at small scales or by generating small-scale cold dark matter isocurvature.
@@ II.1 Enhanced curvature power spectrum
We start by considering primordial blue-tilted curvature power spectra at small scales, which can be produced by e.g. [26, 18, 19, 20, 21, 22, 23, 25, 27, 28, 29, 48, 49, 31, 30]. The primordial spectrum may possess additional features, such as a turnover that produces a bump at some large value of $k$. Since such features are model dependent, we do not account for them. However, we impose a threshold of $\Delta_{\zeta}^{2}(k)\leq 0.01$, forcing the power spectrum to remain constant at arbitrarily large $k$. The presence of the threshold has no significant impact on our analyses. [^1] There are cases in which the threshold is imposed at scales relevant to our analyses. However, these situations occur in the signal-dominated regime, so our projections based on signal-to-noise ratios remain insensitive to the threshold.
To investigate small-scale enhancements in a model-independent manner, we adopt a simple broken-power-law parametrization of the dimensionless primordial curvature power spectrum:
@@
The power-law scaling for $k>k_{b}$ in Eq. (2) can be mapped to various models. For example, for tilted hybrid inflation [27], the spectral index and the amplitude of the power spectrum can be mapped to the effective mass of the inflaton field and the value of the field at the moment when perturbations that have momentum $k$ at the end of inflation were produced. Another example is quantum decoherence during inflation [48]. The evolution of inflationary curvature perturbations can be modified due to the interaction with other degrees of freedom present in the early Universe. This interaction leads to quantum decoherence, and in various cases to a blue-tilted correction of the power spectrum on scales considered in this work. The time dependence and strength of the interaction between the perturbations and other degrees of freedom determine $n_{b}$ and $k_{b}$.
@@ II.2 Enhanced dark matter isocurvature power spectrum
A different class of small-scale enhancements arises from the production of isocurvature modes [34, 35, 94, 43, 62, 95, 47, 24, 36, 38, 41, 39, 40, 32, 37, 44, 33, 46]. We assume that adiabatic modes give rise to the dimensionless primordial curvature power spectrum in Eq. (1), and the isocurvature modes produce a dimensionless primordial isocurvature power spectrum
@@
as demonstrated in Fig. 2. For all scenarios considered in this work, we have checked explicitly that the amplitudes are consistent with constraints from Planck [7]. Similar to the curvature case, we impose a threshold of $\Delta^{2}(k)\leq 0.01$.
@@
A blue-tilted isocurvature power spectrum can arise, for example, from the QCD axion [44, 47, 96, 32], which produces an isocurvature spectral index of $1<n_{\mathrm{iso}}\leq 4$. The bounds on the isocurvature parameters can be translated to a bound on the inflationary energy scale $H_{I}$ [97, 98]. A similar translation can be done in the context of ultralight axions [99], for which $H_{I}$ depends on $A_{\rm iso}/A_{s}$ and the fraction of dark matter that ultralight axions constitute.
Alternatively, a blue tilt with $n_{\rm iso}=4$ can be generated by particle production mechanisms with a finite correlation length. The resulting isocurvature power spectrum vanishes on scales smaller than a cutoff scale determined by causality, and only white noise is left on these smaller scales. Examples of such mechanisms are postinflationary dark matter production [100, 101], vector dark matter [41, 102], and Peccei Quinn breaking during inflation with subsequent axion production [103].
@@ II.3 Running parameters
The parametrization in Eq. (2) is meant to describe the main features of the primordial spectrum; however, if there is a transition from the decreasing to the increasing behavior as a function of $k$, this transition would be described by the first and higher order derivatives of the spectrum. We parametrize the primordial power spectrum in this case as [104]
@@
are the first and second running parameters of the tilt, respectively [104, 105]. Planck 2018 data constrains the running parameters to be $\alpha_{s}=0.013\pm 0.012$ and $\beta_{s}=0.022\pm 0.012$ (at the pivot scale $k=0.05\,\rm{Mpc}^{-1}$[1, 106, 7].
Standard single-field slow-roll inflation predicts $\alpha_{s}\sim(1-n_{s})^{2}\sim\mathcal{O}(\epsilon^{2},\eta^{2})\sim 10^{-3}$ and $\beta_{s}\sim(1-n_{s})^{3}\sim\mathcal{O}(\epsilon^{3},\eta^{3})\sim 10^{-5}$ [107, 89, 104, 108]. Therefore, high precision constraints on the running parameters could either confirm or rule out this inflationary scenario and permit tests of other inflationary models that predict a running [109, 110, 111, 89], such as multifield [112], and warm inflation [113].
Note that this parametrization is only valid for small deviations from the scale-invariant case and is thus not appropriate for the large enhancements studied in Sec. II.1. We focus on the region of $k$ in which the power spectrum begins to transition to becoming blue tilted, and thus the parametrization is valid.
@@
In this section we describe the main observable of interest: the intensity mapping signal of 21-cm neutral hydrogen. We provide a brief review of key concepts for 21-cm LIM from the dark ages, before the formation of the first stars, which occurs around $z\sim 30$ in $\Lambda$CDM. [^2] We note that enhancements of small-scale structure should lead to advanced star formation. Studying the onset of cosmic dawn is beyond the scope of this work, and we use $z=30$ as a convenient reference redshift for our forecasts. We assume no Lyman-$\alpha$ coupling throughout the entire dark ages, including $z=30$. For more details, we refer the reader to Refs. [114, 66, 115, 116, 117] for 21-cm LIM, Refs. [65, 64, 118, 119, 63, 120] for specific application to the dark ages, and Refs. [121, 122] for recent reviews on LIM.
@@
The observed quantity for the 21-cm signal is the brightness temperature contrast with respect to CMB [93],
@@
The Sobolev optical depth [123] is
@@
The brightness temperature is a function of the local hydrogen density and the gas temperature, and it can be parametrized as [93]
@@
where $\delta_{b}=\delta n_{b}/\bar{n}_{b}\simeq\delta_{H}$ up to negligible corrections, $\delta_{T_{\rm gas}}=\delta T_{\rm gas}/\bar{T}_{\rm gas}$, and $\delta_{v}=-(1+z)\partial_{r}v/H(z)$. The coefficients $c_{b,T}$ are functions of redshift only; the mean brightness temperature $\bar{T}_{21}$ is defined by setting all perturbations to zero. Following Ref. [124], we neglect fluctuations in the gas temperature and in the neutral hydrogen fraction to write the 21-cm brightness temperature fluctuations as
@@
By combining Eq. (11) with the continuity equation, we can directly relate the 21-cm angular power spectrum to the matter power spectrum $P_{m}(k)$ [125, 93, 124]:
@@
where $X=\{\zeta,\mathcal{S}_{\rm cdm}\}$ represents the initial condition, and we assume the adiabatic and cold dark matter isocurvature modes are uncorrelated. We use CLASS [126] to compute the linear growth factor $D_{X}(z)$ and the matter transfer function $T_{m,X}(k)$. In the $\Lambda$CDM case, the power spectrum is still linear for the scales and redshifts we probe. In the cases where the primordial power spectrum reaches values above $10^{-3}$ at the scales of interest, the resulting matter power spectrum could become nonlinear. For our results, the majority of the sensitivity is near the $\ell_{\rm{max}}$ value (see Sec. V for more details), and we do not expect our conclusions to be affected by assuming the matter power spectrum is still linear; a careful investigation of this issue is left for future work.